OpenStudy (sydney3545):

Solve 5x2 = -30x - 65. A.) x = -3 ± 2i B.) x = -3 ± 4i C.) x = -6 ± 2i D. x = -6 ± 4i

OpenStudy (danjs):

5x^2 +30x + 65 = 0 A=5 B=30 C=65 \[x = \frac{ -B \pm \sqrt{B^2 - 4*A*C} }{ 2*A }\]

OpenStudy (campbell_st):

rewrite the equation and divide every term by 5 \[x^2 + 6x + 13 = 0\] now use the general quadratic formula

OpenStudy (sydney3545):

I got x=-6 ± 4i not sure if im right

OpenStudy (campbell_st):

well its close just simplify \[x = \frac{-6 \pm 4i}{2}\] remove the common factor

OpenStudy (sydney3545):

So would it be D?

OpenStudy (campbell_st):

well there is a 2 in the denominator, and both terms in the numerator are positive so halve the values in numerator

OpenStudy (campbell_st):

oh, its not D

OpenStudy (sydney3545):

so what would it be haha

OpenStudy (sydney3545):

Wait it would be A

OpenStudy (campbell_st):

correct

OpenStudy (sydney3545):

Solve -3x2 - 4x - 4 = 0. Would this be C? A.) x equals quantity of 2 plus or minus 4i square root of 2 all over 3 B.) x equals quantity of 2 plus or minus 2i square root of 2 all over 3 C.)x equals quantity of negative 2 plus or minus 2i square root of 2 all over 3 x eD.)quals quantity of negative 2 plus or minus 4i square root of 2 all over 3

OpenStudy (campbell_st):

well the general quadratic formula is \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] then multiply every term by -1 so you get, just makes things easy a = 3 b = 4 and c = 4 so the roots are \[x = \frac{-4 \pm \sqrt{16 - 48}}{6}\] now simplify

OpenStudy (sydney3545):

ok im getting C still